OpenStudy (anonymous):

the halfway point of a tunnel through a mountain is 3/2 miles from either end of the tunnel. The mountain is 660 feet (1/8 mile) high. Find the slope of the side of the mountain.

OpenStudy (anonymous):

slope is the tangen of the angle which adjacent to 3/2 milse side and opposite to the 1/8 miles side. Therefore slope = opposite/adjacent = (1/8)/(3/2) = (1/8) / ( 12/8) = 1/12

OpenStudy (anonymous):

thanks